Which of the following statements is/are true?
A. The probability that two numbers are the same is greater than the probability that all the
numbers are different.
B. The probability of not obtaining a 2 is the same as the probability of obtaining only the numbers
1, 3 or 4.
C. The probability of scoring both 1 and 3 is 9
32 .
1. Only A 2. Only B 3. Only C
4. Only A and B 5. A, B and C
1
Expert's answer
2019-08-01T13:28:41-0400

Let it be about three dice rolls with four faces numbered from 1 to 4

A)The probability that all numbers are different: P=4*3*2/64=3/8

The probability that two numbers at two throws coincided: P=(4*4*1/64)*3=3/4 (multiplication by three is due to the fact that the combinations of matches of cubes 3: 1 and 3; 2 and 3 ; 1 and 2;)

3/4>3/8=> A true;

B)This statement is true because there are no other options for not dropping 2 as 1, 3 or 4

C) P=(4*2*1/64)*3 =3/8;(multiplication by three is due to the fact that the combinations of matches of cubes 3: 1 and 3; 2 and 3 ; 1 and 2;)

3/89/36 => C false;

Anwser:4